Anzio Inc. has two classes of shares. Class B has 15 times the voting rights as Class A. If you own 11% of the Class A shares and 20% of the Class B shares, what percentage of the total voting rights do you hold?
Answer #1

Hi

1 2 3 I AB I A 11% 0.066667 B 20% 1 Total voting D 0.73% 20.00% 20.73%

=1/15 1 2 3 A B в 0.11 0.2 Total voting =B1*c1 =B2*C2 =SUM(D1:02)